40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5089     Entry time: Tue Aug 2 02:35:23 2011
Author: kiwamu 
Type: Update 
Category: General 
Subject: preparation of the vent : status and plan 

The vent will take place on Wednesday.

Plan for Tuesday :

  (Morning) Preparation of necessary items for the low power MC (Steve / Jamie)

  (Daytime) Measurement of the MC spot positions (Suresh)

  (Daytime) Arm length measurement (Jenne)

  (Nighttime) Locking of the low power MC (Kiwamu / Volunteers)

 

Plan for Wednesday :

  (Early morning) Final checks on the beam axis, all alignments and green light (Steve / Kiwamu / Volunteers )

  (Morning) Start the vent (Steve)

  (daytime-nighttime) Taking care of the Air/Nitrogen cylinders (Everybody !!)

 

Status of the vent preparation :

 

  (not yet) Low power MC

  (ongoing) Measurement of the arm lengths

  (ongoing) Measurement of the MC spot positions

  (80% done) Estimation of the tolerance of the arm length (#5076)

  (done) Alignment of the Y green beam (#5084)

  (done) Preparation of beam dumps (#5047)

  (done) Health check of shadow sensors and the OSEM damping gain adjustment (#5061)

  (done) Alignment of the incident beam axis (#5073)

  (done) Loss measurement of the arm cavities (#5077)

ELOG V3.1.3-